OpenStudy (anonymous):

Find the derivative of the function. y=(cos(x-1))/(x-1)

OpenStudy (anonymous):

use quotient rule \[\Large \frac{{\frac{d}{dx}(\cos(x-1)*(x-1)-\frac{d}{dx}(x-1)*\cos(x-1)}}{(x-1)^2}\]

OpenStudy (anonymous):

can you do this now ?

OpenStudy (anonymous):

yes, thank you.

OpenStudy (anonymous):

welcome:)